OpenStudy (anonymous):

solving linear equations 2x + 3x = 4+2x+2

OpenStudy (anonymous):

2x + 3x = 4 + 2x + 2 5x = 6 + 2x Subtract 2x from both sides 3x = 6 Divide both sides by 3 x = 2